# Satellite Data Shows Resilience of Tigrayan Farmers in Crop Cultivation During Civil War

### v1

@@ 4.1 Key findings and agreement with other studies
[…]
@@ 4.2 Limitations
We followed best practices established by the scientific literature whenever possible, but there are some limitations inherent in our approach. One possible source of error in our estimates is in the reference labels used to estimate area. We tried to maximize the correctness and confidence of these reference labels as much as possible through our response design. We used high-resolution images from PlanetScope (3 m/pixel) and Google Earth ($<$1 m/pixel) when available. Each sample was analyzed and labeled independently by at least two different annotators; if there was any disagreement between annotators, we convened expert annotators to review and discuss all of the available evidence and come to a consensus label for the sample. Even with these precautions, it is possible that some samples were interpreted incorrectly due to the difficulty of differentiating between crop fields that are fallow and fields that are planted from satellite images, even in high-resolution (3 m/pixel) images (see Figure 3).
@@
Another limitation of this study is that we used only two years for our change analysis: 2021 (mid-war) and 2020 (pre-war). Including additional pre-war years in the analysis would help establish a baseline of “nominal” area estimates to compare with mid-war 2021. It is also possible that in addition to the outbreak of war during the end of the harvest period in 2020, the COVID-19 pandemic may have had some impact on the 2020 growing season [44]. Nevertheless, we included only 2020 and 2021 in our study due to the significant resources required for analyzing two years and the precedence set by prior studies using similar statistical techniques that compared two years for change assessments (e.g., [20, 7, 11]).
@@
Finally, the impacts of the Tigray war on crop cultivation in our study were limited to those that resulted in a binary change in cultivation status: either cultivated or not cultivated. This study did not account for other impacts such as population displacement, changes in crop types or cultivation practices, movement and access limitations, market dysfunction and loss of income, or destruction of food stocks and farming equipment that were reported in farmer surveys [6, 3]. While some studies have analyzed ploughing status and harvest status from satellite observations [3, 19], most of the other factors listed are difficult or impossible to determine from satellite observations, especially in the absence of extensive ground-truth data for calibration.
@@ 5 Conclusions
Despite widespread disruptions to agriculture, our remote sensing-based crop maps and area estimates suggest Tigrayan smallholder farmers were largely able to carry out crop cultivation activities in 2021 amidst the Tigray war. Our approach using satellite remote sensing, machine learning, and statistical techniques enables timely, accurate insights about agriculture in conflict regions where ground access is not possible using satellite data and statistical methods. Further studies could build on our methodology to incorporate additional years before and during the Tigray war to continue monitoring Tigray’s agriculture and food security situation. We made all code and data used in this study publicly accessible to enable transparency, reproducibility, and future work.
@@ Acknowledgments
This work was supported by NASA under Award #80NSSC21K1511 and the NASA Harvest Consortium on Food Security and Agriculture (Award #80NSSC18M0039). We thank the Famine Early Warning Systems Network (FEWS NET) Science and Early Warning Teams for assistance with point labeling and discussions on early versions of results. We thank Liya Weldegebriel, Emnet Negash, and Jan Nyssen for sharing their concurrent findings, local expertise, and ground-truth data which helped strengthen this study.
Appendix A Sentinel-2 time series of cultivated and fallow fields
@@
Appendix B Admin2-scale confusion matrices
@@
Appendix C Admin2-scale change map accuracy
@@
Table 7: Accuracy metrics for four-class cropland change reference sample in Southern zone of Tigray. Overall accuracy is $0.65\pm 0.04$.
@@
Table 8: Accuracy metrics for four-class cropland change reference sample in Central zone of Tigray. Overall accuracy is $0.73\pm 0.05$.
@@
Table 9: Accuracy metrics for four-class cropland change reference sample in Western zone of Tigray. Overall accuracy is $0.66\pm 0.06$.
@@
Table 10: Accuracy metrics for four-class cropland change reference sample in Eastern zone of Tigray. Overall accuracy is $0.74\pm 0.06$.
@@
Table 11: Accuracy metrics for four-class cropland change reference sample in North Western zone of Tigray. Overall accuracy is $0.71\pm 0.05$.
Appendix D Conflict buffer confusion matrices
@@
Appendix E Conflict buffer change map accuracy
@@
Table 12: Accuracy metrics for four-class cropland change reference sample inside 5km buffer of conflict events in Tigray. Overall accuracy is $0.78\pm 0.05$.
@@
Table 13: Accuracy metrics for four-class cropland change reference sample outside 5km buffer of conflict events in Tigray. Overall accuracy is $0.77\pm 0.03$.
qWlUonKsvwWwD8ymc/nXwVBeLjf7xEKhdBut9Hr9WgmkyGEkJwsy5eHG5vN5g0AKIoCAEgkEkin0wQAfN9/cXPdheu6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ 3.4 Cropland area changes in proximity to conflict events
Since the reference sample size for the area inside the conflict buffer ($n=219$) was smaller than the sample size for the area outside the conflict buffer ($n=582$), the higher upper bound of loss in the conflict buffer could be due to a relatively smaller sample size. To test this, we performed an experiment to estimate the cropland loss area outside the conflict buffer with a smaller sample size equal to that inside the conflict buffer. We randomly sub-sampled 219 points from the non-conflict 582 reference samples, then calculated the cropland loss area and confidence interval (CI) using this sample. We repeated this 10 times using different random seeds, since the result may vary depending on the samples drawn. We calculated the median and mean cropland loss area and CI over all random seeds. Table 14 reports the results from this experiment and Table 15 summarizes the results compared to those reported in Table 6 using the full reference sample. For both the mean and median estimates with $n=219$, we found a lower upper bound CI as a percentage of total area ($0-1\%$ for the median, $0-2\%$ for the mean), consistent with our result using the larger sample size ($0-1\%$). We note that the mean estimate is sensitive to outliers thus we feel the median is a more representative estimate. These results confirm that the higher upper bound of cropland loss inside the 5 km conflict buffer is most likely due to a higher likelihood of loss near conflict events, rather than a smaller sample size.
@@ 4.1 Key findings and agreement with other studies
In reports published in the first half of 2021 prior to the planting and growing season, there were grave concerns and uncertainty about whether Tigrayan farmers would be able to cultivate their crops in the 2021 meher season [18, 3]. Satellite remote sensing data provides real-time and historical observations that can be used for qualitative and quantitative assessments of how conflicts affect agricultural land use and production. This source of data is particularly valuable for evaluating impacts during armed conflict, since traditional ground-based surveys may be impractical due to safety concerns. Using remote sensing observations of Tigray during and prior to the war, studies such as ours, Nyssen et al. [3], and Weldegebriel et al. [19] could ascertain information to help clarify uncertainty about the impact of the conflict on cultivated area.
Our finding that crop cultivation area in Tigray in wartime 2021 was comparable to cultivated area in pre-war 2020 agrees with previous studies of the impact of the war on cultivated area [3, 19]. Through ground interviews and satellite image analysis, these studies concluded that despite many reports of disruption to farming activities, most Tigrayan farmers were ultimately able to plant and harvest crops by employing wartime coping strategies such as shifting the timing and location of farming activities, changing crop types, and leveraging communal aid [3]. While we did not find a net loss of cropland as in Weldegebriel et al. [19], it is difficult to directly compare the 714 km2 net loss found by their study to our study since their study was restricted to highland croplands in Tigray, employed a specific definition of “well-cultivated cropland,” estimated area using pixel counting instead of a reference sample, and did not provide confidence intervals. Nevertheless, our estimate of area attributed to cropland loss in Tigray was between 19-79 kha which is comparable to the net loss of 714 km2 (or 71.4 kha) found by Weldegebriel et al. [19].
@@ 4.2 Limitations
We followed best practices established by the scientific literature whenever possible, but there are some limitations inherent in our approach. One possible source of error in our estimates is in the reference labels used to estimate area. We tried to maximize the correctness and confidence of these reference labels as much as possible through our response design. We used high-resolution images from PlanetScope (3 m/pixel) and Google Earth ($<$1 m/pixel) when available. Each sample was analyzed and labeled independently by at least two different annotators. If there was any disagreement between annotators, we convened expert annotators to review and discuss all of the available evidence and come to a consensus label for the sample. Even with these precautions, it is possible that some samples were interpreted incorrectly due to the difficulty of differentiating between crop fields that are fallow and fields that are planted from satellite images, even in high-resolution (3 m/pixel) images (see Figure 3). A shows some examples of reference sample labels and corresponding high-resolution satellite images from the reference dataset.
@@
Another limitation of this study is that we used only two years for our change analysis: 2021 (wartime) and 2020 (pre-war). Including additional pre-war years in the analysis would help establish a baseline of “nominal” area estimates to compare with wartime 2021. It is also possible that in addition to the outbreak of war during the end of the harvest period in 2020, the COVID-19 pandemic may have had some impact on the 2020 growing season [47]. Nevertheless, we included only 2020 and 2021 in our study due to the significant resources required for analyzing two years and the precedence set by prior studies using similar statistical techniques that compared two years for change assessments (e.g., [20, 6, 10]).
Finally, the impacts of the Tigray war on crop cultivation in our study were limited to those that resulted in a binary change in cultivation status: either cultivated or not cultivated. We designed our study this way because our primary goal was to provide early warning insights to the USAID Famine Early Warning Systems Network’s Early Warning Team about the potential for shortfalls in crop production in Tigray during the war. Our goal was to broadly assess whether the total cropland area under cultivation in Tigray was lower during the war (2021) compared to prior years (2020). Thus, we followed a similar approach to previous work that estimated the change in cropland area between a pre-war and wartime period (e.g., [6, 9, 7, 15, 8]).
@@
Our study did not account for other impacts such as population displacement, changes in crop types or cultivation practices, movement and access limitations, market dysfunction and loss of income, or destruction of food stocks and farming equipment that were reported in farmer surveys [18, 3]. While some studies analyzed plowing status and harvest status from satellite observations [3, 19], most of the other factors listed are difficult or impossible to determine from satellite observations, especially in the absence of extensive ground-truth data for calibration. Other studies used quasi-experimental methods such as propensity score matching to account for additional covariates such as population density, distance to roads, or distance to conflict events [10]. We did not employ such techniques since the objective of our study was a broad assessment of cropland area change occurring between the pre-war (2020) and wartime (2021) period. In addition, as shown in Figure 1(a), conflict events during the study period were widely distributed and not aggregated along a front or in clusters. Using the distance to conflict or the density of conflicts as a predictor variable may not adequately capture the nuanced impacts that individual conflict events can have on cropland loss due to the heterogeneous nature of warfare tactics, varying levels of intensity, and the different scales at which these events occur.
@@ 5 Conclusions
Despite widespread disruptions to agriculture, our remote sensing-based crop maps and area estimates suggest Tigrayan smallholder farmers were largely able to carry out crop cultivation activities in 2021 amidst the Tigray war. Our approach using satellite remote sensing, machine learning, and statistical techniques enables timely, accurate insights about agriculture in conflict regions where ground access is not possible using satellite data and statistical methods. Further studies could build on our methodology to incorporate additional years before and during the Tigray war to continue monitoring Tigray’s agriculture and food security situation. We made all code and data used in this study publicly accessible to enable transparency and reproducibility and to facilitate future work.
@@ Acknowledgments
This work was supported by NASA under Award #80NSSC21K1511 and the NASA Harvest Consortium on Food Security and Agriculture (Award #80NSSC18M0039). We thank the Famine Early Warning Systems Network (FEWS NET) Science and Early Warning Teams for assistance with point labeling and discussions on early versions of results. We thank Liya Weldegebriel, Emnet Negash, and Jan Nyssen for sharing their concurrent findings, local expertise, and ground-truth data which helped strengthen this study. Finally, we thank the anonymous journal reviewers for their constructive feedback that greatly improved this study.
Appendix A Examples of reference sample interpretations from satellite images
In this section, we provide several examples of reference data samples to illustrate how we interpreted change class labels for each point based on high-resolution satellite images from Planet Monthly Mosaics and Google Earth points. In the caption of each figure, we provide the ID of the point in change_2020-2021_strat_ref_samples_labeled.shp file provided at https://data.harvestportal.org/dataset/annual-and-change-maps-for-tigray-2020-2021. We indicate the approximate location of each point within Tigray by the gold diamond on each figure. Red dots indicate the point was labeled as non-crop/not cultivated while green dots indicate the point was labeled as crop/cultivated.
Figure 8: Example point labeled as planted cropland loss (id #238). The reference point is in a field that appears cultivated in 2020 but fallow in 2021, although there are adjacent fields that are cultivated in 2021. The first column shows images from Airbus/CNES from Google Earth while other columns show Planet Monthly mosaics from Collect Earth Online.
Figure 9: Example point labeled as planted cropland loss (id #485). The reference point is in a field that appears cultivated in 2020 but fallow in 2021. In June 2020, the field containing the point appears dark brown suggesting land preparation; in September, it appears dark green and uniformly vegetated; in November, it appears to have been cleared/harvested. However, in 2021, nearby fields show signs of preparation with dark brown soil in June but not the reference field; in September, the field has some faint greenness but is sparse and patchy; in November, the field appears to have sparse and lingering vegetation indicative of a fallow field. The first column shows an image from Airbus/CNES from Google Earth (no high-resolution image was available for 2020) while other columns show Planet Monthly mosaics from Collect Earth Online. The Google Earth image was taken outside of the growing season, but still provides helpful context showing the area in higher resolution than can be seen in the Planet mosaics.
Figure 10: Example point labeled as planted cropland gain (id #549). The reference point is in a field that appears fallow in 2020 but cultivated in 2021. In 2020, the field containing the point does not exhibit signs of land preparation, as it remains light brown in June 2020, or of vegetative growth, as it remains brownish/dull green in September 2020. In June 2021, the field and other fields surrounding it appear to have a darker brown color suggesting land preparation; in September, it appears bright green and uniformly vegetated; in November, it appears to have been cleared/harvested. The first column shows images from Airbus/CNES (2020) and Maxar (2021) from Google Earth while other columns show Planet Monthly mosaics from Collect Earth Online. The Google Earth images were taken outside of the growing season, but still provide helpful context showing the area in higher resolution than can be seen in the Planet mosaics.
Figure 11: Example point labeled as planted cropland gain (id #523). The reference point is in a field that appears fallow in 2020 but cultivated in 2021, which is clearly shown in the high-resolution Google Earth images from October 2020 versus 2021 shown in the second column. In August 2020, the field appears dull green compared to the brighter, uniform green shown in the September 2021 Planet mosaic. The first two columns show images from Maxar from Google Earth while other columns show Planet Monthly mosaics from Collect Earth Online.
Figure 12: Example point labeled as planted cropland gain (id #206). The reference point is in a field that appears fallow in 2020 but cultivated in 2021. The Google Earth image from June 2021 clearly shows that the field has been prepared for planting, which is also seen in the July 2021 Planet mosaic but not in 2020. The field appears densely green and vegetation in August 2021 but bare in the September 2020 mosaic. The first column shows an image from Maxar from Google Earth (no high-resolution image was available for 2020) while other columns show Planet Monthly mosaics from Collect Earth Online.
Figure 13: Example point labeled as stable non-crop (id #0). The reference point is in a terraced area that could have been used for agriculture in prior years, but in 2020 and 2021 appears uncultivated. This can be seen in the high-resolution Google Earth image from November 2021 which shows patchy vegetation and the Planet Monthly mosaics which do not show signs of land preparation, dense vegetation during the growing period, or harvesting.
Appendix B Sentinel-2 time series of cultivated and fallow fields
@@
Appendix C Zone-scale confusion matrices
@@
Appendix D Zone-scale change map accuracy
@@
Table 7: Accuracy metrics for four-class cropland change reference sample in Southern zone of Tigray. Overall accuracy is $0.65\pm 0.04$. Legend: UA = User’s Accuracy, PA = Producer’s Accuracy.
@@
Table 8: Accuracy metrics for four-class cropland change reference sample in Central zone of Tigray. Overall accuracy is $0.73\pm 0.05$. Legend: UA = User’s Accuracy, PA = Producer’s Accuracy.
@@
Table 9: Accuracy metrics for four-class cropland change reference sample in Western zone of Tigray. Overall accuracy is $0.66\pm 0.06$. Legend: UA = User’s Accuracy, PA = Producer’s Accuracy.
@@
Table 10: Accuracy metrics for four-class cropland change reference sample in Eastern zone of Tigray. Overall accuracy is $0.74\pm 0.06$. Legend: UA = User’s Accuracy, PA = Producer’s Accuracy.
@@
Table 11: Accuracy metrics for four-class cropland change reference sample in North Western zone of Tigray. Overall accuracy is $0.71\pm 0.05$. Legend: UA = User’s Accuracy, PA = Producer’s Accuracy.
Appendix E Conflict buffer confusion matrices
@@
Appendix F Conflict buffer change map accuracy
@@
Table 12: Accuracy metrics for four-class cropland change reference sample inside 5km buffer of conflict events in Tigray. Overall accuracy is $0.78\pm 0.05$. Legend: UA = User’s Accuracy, PA = Producer’s Accuracy.
@@
Table 13: Accuracy metrics for four-class cropland change reference sample outside 5km buffer of conflict events in Tigray. Overall accuracy is $0.77\pm 0.03$. Legend: UA = User’s Accuracy, PA = Producer’s Accuracy.
Appendix G Effect of sample size on conflict buffer estimates
Since the reference sample size for the area inside the conflict buffer ($n=219$) was smaller than the sample size for the area outside the conflict buffer ($n=582$), it is possible that the higher upper bound of loss in the conflict buffer reported in Table 6 is due to a relatively smaller sample size. To test this, we performed an experiment to estimate the cropland loss area outside the conflict buffer with a smaller sample size equal to that of the conflict buffer. We randomly sub-sampled 219 points from the non-conflict 582 reference samples, then calculated the cropland loss area and CI using this sample. We repeated this 10 times using different random seeds, since the result may vary depending on the samples drawn. We calculated the median and mean cropland loss area and CI over all random seeds. We report the results in Table 14.
Seed Cropland loss area (kha) OA UA PA
1 $26\pm 28$ 0.74 $\pm$ 0.05 0.10 $\pm$ 0.11 1.00 $\pm$ 0.00
10 $54\pm 61$ 0.79 $\pm$ 0.05 0.10 $\pm$ 0.09 0.47 $\pm$ 0.54
100 $22\pm 43$ 0.77 $\pm$ 0.05 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
1000 $21\pm 23$ 0.76 $\pm$ 0.05 0.08 $\pm$ 0.09 1.00 $\pm$ 0.00
10000 $41\pm 48$ 0.77 $\pm$ 0.05 0.08 $\pm$ 0.08 0.47 $\pm$ 0.56
100000 $33\pm 27$ 0.77 $\pm$ 0.05 0.13 $\pm$ 0.11 1.00 $\pm$ 0.00
2 $15\pm 20$ 0.75 $\pm$ 0.05 0.06 $\pm$ 0.08 1.00 $\pm$ 0.00
20 $28\pm 27$ 0.76 $\pm$ 0.05 0.11 $\pm$ 0.10 1.00 $\pm$ 0.00
200 $51\pm 70$ 0.73 $\pm$ 0.05 0.07 $\pm$ 0.07 0.33 $\pm$ 0.50
2000 $29\pm 27$ 0.78 $\pm$ 0.05 0.11 $\pm$ 0.11 1.00 $\pm$ 0.00
Median $29\pm 28$
Mean $32\pm 37$
Table 14: Cropland loss area, overall accuracy (OA), user’s accuracy (UA), and producer’s accuracy (PA) for random subsets of size $n=219$ from reference sample in region outside of the 5 km conflict buffer.
Cropland loss area (kha) Percentage of total area
Inside buffer (n=219) 19 $\pm$ 17 0–3%
Outside buffer (n=582) 28 $\pm$ 23 0–1%
Outside buffer (n=219, median) 29 $\pm$ 28 0–1%
Outside buffer (n=219, mean) 32 $\pm$ 37 0–2%
Table 15: Comparison of cropland loss area estimates inside and outside the 5 km conflict buffer for full sample sizes (reported in Table 6) and equal sample sizes.